# Edge-guided Low-light Image Enhancement with Inertial Bregman Alternating Linearized Minimization

### v1

@@ Appendix A Proof of Lemma 1
This together the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (24) and $\delta=\frac{\varepsilon}{2}\min\{\delta_{1},\delta_{2}\}$, we can obtain (25) easily. This completes the proof. ∎
Appendix B Proof of Lemma 2
@@
$0\in$ $\partial\theta_{1}(x_{1}^{k+1})+\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k})-\nabla_{x_{1}}h^{-}(z_{1}^{k},x_{2}^{k})$ (34)
@@
$0\in$ $\partial\theta_{2}(x_{2}^{k+1})+\nabla_{x_{2}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},z_{2}^{k})$ (35)
@@
First, from the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (24), we have
@@
where the last inclusion follows from (34). Second, we compute the subgradient of $\Theta_{\delta_{1},\delta_{2}}$ with respect to $w_{12}$,
@@
where the inclusion follows from (35). Finally, it follows from (24) that
@@
$\quad\lambda_{2}^{+}\|x_{2}^{k}-x_{2}^{k+1}\|+(\lambda_{2}^{+}\bar{\beta}_{2}+\tau_{2}^{+}\bar{\alpha}_{2})\|x_{2}^{k}-x_{2}^{k-1}\|$
@@
Appendix C Proof of Theorem 1
@@
We first prove (i). It follow from Lemma 1 that for any $N\in\mathbb{N}$,
@@
This fact together with (25) implies that $\left\|\mathrm{w}^{k}\right\|\rightarrow 0$ as $k\rightarrow\infty$.
@@
Consequently, $\lim_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}\left({\bf w}^{k_{j}}\right)=\Theta^{*}$. On the other hand, it follows from Lemmas 2 that
@@
Now we prove (iii). According to Assumption 1, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is bounded below. Besides, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is nonincreasing from (25). Thus, the limit $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta^{\infty}$ exists. Taking any ${\bf w}^{*}\in z({\bf w}^{0})$ as an accumulation point of $\{{\bf w}^{k}\}$, then there exists a subsequence $\{{\bf w}^{k_{j}}\}$ satisfying $\lim_{j\rightarrow\infty}{\bf w}^{k_{j}}={\bf w}^{*}$. From Items (i) and (ii), we have
@@
Thus, we complete the proof by the arbitrariness of ${\bf w}^{*}$ in $z({\bf w}^{0})$. ∎
Appendix D Proof of Theorem 2
@@
Thus, we have ${\bf w}^{k+1}={\bf w}^{k}$ for any $k>\bar{k}$. Thus the assertion $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty$ holds trivially. Otherwise, since $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})$ is nonincreasing from Lemma 1, we have $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})>\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$ for all $k$. Again from $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$, we know that for any $\eta>0$, there exists a nonnegative integer $k_{0}$ such that $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})<\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})+\eta$ for any $k>k_{0}$. In addition, for any $\varsigma>0$ there exists a positive integer $k_{1}$ such that ${\rm dist}({\bf w}^{k},z({\bf w}^{0}))<\varsigma$ for all $k>k_{1}$. Consequently, for any $\eta,~{}\varsigma>0$, when $k>k_{2}:=\max\{k_{0},k_{1}\}$, we have
@@
$\zeta_{p,q}:=$ ${\varphi_{k}}\big{(}\Theta_{\delta_{1},\delta_{2}}(u^{p})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})\big{)}$
@@
[…]
$\begin{split}&\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\\ &\leq\frac{\zeta_{k,k+1}}{{\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))}\\ &\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1}.\end{split}$
@@
$\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}$
$\leq\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})$
$\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1},$
@@
$\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|$
$\leq\sqrt{\frac{\gamma}{\delta}\zeta_{k,k+1}\|{\bf w}^{k}-{\bf w}^{k-1}\|}$

### v2

@@ IV Convergence results
This together the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (22) and $\delta=\frac{\varepsilon}{2}\min\{\delta_{1},\delta_{2}\}$, we obtain (23) and the proof is completed. ∎
Lemma 3.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by Algorithm 1, for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, for any $k\geq 1$, there exists $\gamma>0$ such that
@@
0 0$\in$0 ∈ $\partial\theta_{1}(x_{1}^{k+1})+\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k})-\nabla_{x_{1}}h^{-}(z_{1}^{k},x_{2}^{k})$ (34)
@@
0 0$\in$0 ∈ $\partial\theta_{2}(x_{2}^{k+1})+\nabla_{x_{2}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},z_{2}^{k})$ (35)
@@
Firstly, from the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (22), we have
@@
where the last inclusion follows from (34). Secondly, we compute the subgradient of $\Theta_{\delta_{1},\delta_{2}}$ with respect to $w_{12}$,
@@
where the inclusion follows from (35). By (22), we obtain
@@
$\quad+\lambda_{2}^{+}\|x_{2}^{k}-x_{2}^{k+1}\|+(\lambda_{2}^{+}\bar{\beta}_{2}+\tau_{2}^{+}\bar{\alpha}_{2})\|x_{2}^{k}-x_{2}^{k-1}\|$
@@
Theorem 1.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by Algorithm 1, for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. Then,
  • (i)
  • (ii)
  • (iii)
@@
We first prove (i). From Lemma 1, for any $N\in\mathbb{N}$,
@@
This fact together with (23) implies that $\left\|\mathbf{w}^{k}\right\|\rightarrow 0$ as $k\rightarrow\infty$.
@@
Consequently, $\lim_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}\left({\bf w}^{k_{j}}\right)=\Theta^{*}$. On the other hand, it follows from Lemma 3 that
@@
Now we prove (iii). The sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is bounded below by Assumption 1. The sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is nonincreasing from (23). Thus, the limit $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta^{\infty}$ exists. Taking ${\bf w}^{*}\in z({\bf w}^{0})$ as an accumulation point of $\{{\bf w}^{k}\}$, then there exists a subsequence $\{{\bf w}^{k_{j}}\}$ satisfying $\lim_{j\rightarrow\infty}{\bf w}^{k_{j}}={\bf w}^{*}$. From Items (i) and (ii), we have
@@
We complete the proof by the arbitrariness of ${\bf w}^{*}$ in $z({\bf w}^{0})$. ∎
Theorem 2.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by Algorithm 1, for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. If $\Phi$ in (11) is a KL function. Then, the sequence $\left\{{\bf w}^{k}\right\}_{k\geq 1}$ has finite length, that is, $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty.$ Hence, the whole sequence $\{{\bf w}^{k}\}_{k\geq 1}$ is convergent.
@@
Hence, we have ${\bf w}^{k+1}={\bf w}^{k}$ for any $k>\bar{k}$. Thus the assertion $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty$ holds trivially. Otherwise, since $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})$ is nonincreasing from Lemma 1, we have $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})>\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$ for all $k$. Again from $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$, we know that for any $\eta>0$, there exists a nonnegative integer $k_{0}$ such that $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})<\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})+\eta$ for any $k>k_{0}$. In addition, for any $\varsigma>0$ there exists a positive integer $k_{1}$ such that ${\rm dist}({\bf w}^{k},z({\bf w}^{0}))<\varsigma$ for all $k>k_{1}$. Consequently, for any $\eta,~\varsigma>0$, when $k>k_{2}:=\max\{k_{0},k_{1}\}$, we have
@@
$\zeta_{p,q}=$ ${\varphi_{k}}\big{(}\Theta_{\delta_{1},\delta_{2}}(u^{p})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})\big{)}$
@@
Then, associated with ${\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))>0$, ${\rm dist}(0,\partial\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k}))\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|$ in Lemma 3, and (45), we get
$\begin{split}&\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\leq\frac{\zeta_{k,k+1}}{{\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))}\\ &\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1}.\end{split}$
@@
$\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1}.$
@@
$\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|\leq\sqrt{\frac{\gamma}{\delta}\zeta_{k,k+1}\|{\bf w}^{k}-{\bf w}^{k-1}\|}$
@@
Figure 3: Visual comparison of different edge extraction methods in non-reference Fusion dataset [19]. The first column is low-light images; the second column is our enhanced results; the third and the last columns are the edges obtained by using Eq. (4) and the learned edge by our strategy, respectively.
(a) (a) Input (9.13/36.84)
(b) (b) SMG (19.61/81.57)
(c) (c) Ours (22.94/87.68)
(d) (d) GT
(e) (e) edge of (a)
(f) (f) SMG
(g) (g) Ours
(h) (h) GT
(i) (a) Input (12.77/67.42)
(j) (b) SMG (18.30/83.93)
(k) (c) Ours (21.12/91.16)
(l) (d) GT
(m) (e) edge of (a)
(n) (f) SMG
(o) (g) Ours
(p) (d) GT
Figure 4: Image enhancement results with index PSNR (dB) and SSIM (%). (a) is the input low-light image from dataset Fivek [50]; (b) and (c) are enhancement results by SMG [4] and ours; (d) is the ground truth; (e) is the edge of (a) extracted by the Laplace operator; (f) and (g) are corresponding edges generated by SMG [4] and ours; (h) is the ground truth.
V Numerical experiments
To evaluate the effectiveness of our proposed scheme, we first analyze the proposed edge extractor using some ablation studies. Secondly, the parameters of the proposed model and algorithm are studied. Then, the synthesis and real-world low-light datasets are tested to illustrate the robustness of our method. Moreover, some state-of-the-art methods are compared to demonstrate the superiority of the proposed method.
V-A Implement
In this work, we apply Eq. (4) on 485 training images of the LOL dataset [19] to obtain their corresponding edges. As illustrated in Fig. 2, we extract the edge label from the high-contrast image. Our edge extractor is implemented using the PyTorch framework, and we update it using the Adam optimizer. The initial learning rate is set to $10^{-4}$ and halved every 200200200 epochs. Once the edge network is learned, we integrate it into the proposed framework to enhance the performance of the image Retinex model.
Figure 5: Study on the number of multi-scale residual block (MSRB) modules. Considering the cost and the performance, we set the number of modules as 666 in our experiment.
Figure 6: Comparison of the AMA, ADMM, the proposed scheme, and the proposed scheme with anisotropic TV (6), referred to as Ours_anis.
V-B Effectiveness of edge extractor
[…]
Figure 7: PSNR and NIQE results along with model parameters $\alpha$ and $\beta$.
V-C Algorithm and model analysis
In this subsection, we analyze the efficiency of the proposed model (10) with ROF regularizer (5), i.e., $\Psi(l)=\|\nabla l\|_{1}$, referred to as Ours. First of all, we compare the proposed algorithm with the alternating minimization algorithm (AMA) and ADMM. The energy and PSNR curves presented in Fig. 6 illustrate that with the same model parameter setting, the proposed algorithm converges faster and more accurately. Moreover, we also compare our model with the anisotropic TV (6), referred to as Ours_anis. This implies that the proposed algorithm can handle different regularizers robustly.
Secondly, the algorithm parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, and $\tau_{1}$, $\tau_{2}$ are studied. According to the Assumption 3, $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$ with $0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ and $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}\in(0,1)$, $i=1,2$. In experiments, we set strongly convex modulus $\rho_{1}$ and $\rho_{2}$ to be 111, then we have $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$.
In the experiments, by setting $\bar{\beta}_{i}=1-\epsilon$, $\bar{\alpha}_{i}=\frac{1}{2}-\epsilon$, $\rho_{i}=1$, and $\lambda_{i}^{+}=\beta$, we get $\tau_{i}=\frac{\frac{1+\epsilon}{\epsilon}(2.5+\beta_{i}-2\epsilon)\beta}{1-\alpha_{i}}$. Since $\alpha_{i}$ and $\beta_{i}$ are extrapolation parameters, we first fix them as 0 to find $\tau_{i}$. As analyzed before, $\tau_{i}$ can be decided by $\alpha_{i}$, $\beta_{i}$, and $\beta$. The remaining algorithm parameters are $\tau_{i}$, $i=1,2$, and decided by model parameter $\beta$. Hence, we first study the model parameters $\alpha$ and $\beta$ with fixed $\alpha_{i}$ and $\beta_{i}$. After the optimized $\alpha$ and $\beta$ are obtained, we will focus on the algorithm parameters $\alpha_{i}$ and $\beta_{i}$. We first study $\alpha$ and $\beta$ in interval [1,200][1,200][ 1 , 200 ] with stepsize 101010. We find that our model is quite robust with parameter $\beta$ while the curve is turbulent alone $\alpha$. Hence we narrow the interval to [5,50][5,50][ 5 , 50 ] and [0.1,10][0.1,10][ 0.1 , 10 ] use the linespace to find 202020 numbers, respectively. Fig. 7 plots the PSNR and NIQE results along the model parameters $\alpha\in[5,50]$ and $\beta\in[0.1,10]$. The surface indicates that the proposed model is stable to parameters $\alpha$ and $\beta$ in the given interval. Since curves become quite flat as the parameter $\alpha$ from 333 to 555 along the $\alpha$ axis, which implies choosing $\alpha$ in the interval [10,40][10,40][ 10 , 40 ] ensures the best result. Similarly, the interval of model parameter $\beta$ is chosen at [0.1,5][0.1,5][ 0.1 , 5 ].
With the obtained optimized model parameter $\beta$, the algorithm parameter $\tau_{i}$ then controlled by $\alpha_{i}$ and $\beta_{i}$, $i=1,2$. Hence, only $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$ are studied in the algorithm parameter. For these four algorithm parameters, we choose to tune one parameter in its interval while others are fixed as 0. More specifically, we set the model parameters $\alpha_{i}\in[0.01:0.1:0.5]$ and $\beta_{i}\in[0.01:0.1:1]$. From Fig. 8 and Fig. 9, we know that extrapolation parameters speed up the whole algorithm with convergence.
TABLE II: Quantitative results with PSNR (dB), SSIM (%), NIQE, FADE, and SSEQ values on datasets Fivek, LSRW, and SYD. The BEST result is in red whereas the SECOND BEST one is in blue.
 Dataset  Index   Input RetinexNet [19]  GLADNet [51] URetinex [10]   EFINet [52]     RQ [53]    SMG [4]   Bread [54]    Ours
Fivek PSNR 13.42 12.59 16.46 13.22 17.24 14.79 16.31 15.94 17.82
SSIM 55.24 59.96 65.98 62.18 68.27 55.10 65.14 65.28 70.33
NIQE 3.832 3.860 3.450 3.700 3.655 3.405 7.704 4.097 3.761
FADE 1.922 1.731 2.275 2.335 1.922 1.762 2.023 2.105 1.799
SSEQ 26.28 25.16 27.03 28.11 26.26 35.41 21.38 39.26 19.95
LSRW PSNR 9.16 15.48 16.64 15.99 15.29 10.64 16.47 16.34 16.74
SSIM 19.09 48.61 52.92 52.50 53.69 29.44 58.20 59.60 62.72
NIQE 3.278 4.312 3.364 3.804 3.035 3.531 6.702 3.443 3.142
FADE 1.636 1.618 1.974 2.251 1.929 2.271 1.802 2.435 1.432
SSEQ 27.22 27.74 17.80 19.54 30.96 31.09 13.05 33.79 16.17
SYD PSNR 9.13 14.42 16.72 16.57 14.78 16.64 16.18 14.30 16.82
SSIM 30.33 68.41 70.50 69.97 64.74 70.72 68.12 68.00 66.06
NIQE 4.503 6.211 4.245 4.252 4.043 4.121 5.348 4.128 5.209
FADE 0.951 0.894 1.072 0.961 1.200 1.055 1.243 1.172 0.873
SSEQ 15.45 20.181 15.39 16.91 16.41 23.09 12.68 20.91 15.36
V-D Image enhancement
This subsection compares our method with some state-of-the-art image enhancement models, including the traditional methods (LR3M [14], QSPE [55], DRE [6], RMR [18]) and the deep learning methods (RetinexNet [19], GLADNet [51], URetinex [10], EFINet [52], RQ [53], SMG [4], Bread [54]). Note that LR3M, DRE, RMR, RetinexNet, and URetinex are the methods under the Retinex framework. In particular, SMG introduces an end-to-end neural network that leverages edge information to assist with low-light image enhancement.
TABLE III: Quantitative results with PSNR (dB) and SSIM (%) values on dataset Realblur15. The BEST result is in red whereas the SECOND BEST one is in blue.
Image Index  Input RetiNet URetinex  EFINet  RQ   LR3M   QSPE   DRE   RWR   Ours
Img1 PSNR 8.91 16.93 19.69 16.60 16.56 15.70 12.53 16.35 19.92 20.36
SSIM 24.66 70.78 86.65 83.72 79.48 79.26 67.27 77.80 84.98 87.23
Img2 PSNR 8.47 14.41 25.16 15.36 13.12 14.33 11.52 15.35 19.13 18.61
SSIM 29.65 68.72 87.29 79.33 64.81 73.68 68.03 70.98 83.52 85.96
Img3 PSNR 9.49 16.82 23.76 16.34 15.36 15.98 13.22 16.75 17.85 20.33
SSIM 23.97 68.97 88.83 80.38 76.21 79.34 67.77 82.22 80.16 87.02
Img4 PSNR 13.09 15.07 13.63 21.76 13.56 19.19 15.83 13.99 15.85 19.26
SSIM 23.90 44.98 58.00 82.39 57.15 72.36 53.78 51.71 59.85 68.01
Img5 PSNR 10.16 15.25 17.13 16.14 17.16 18.80 13.29 14.83 14.14 19.01
SSIM 23.81 72.64 81.59 86.83 81.58 87.60 61.29 81.71 75.39 88.26
Img6 PSNR 6.71 14.01 21.61 14.82 10.72 12.20 9.22 11.71 15.05 15.98
SSIM 34.67 64.51 88.86 74.78 63.95 70.66 62.39 66.63 83.54 83.07
Img7 PSNR 10.44 15.28 15.56 19.54 25.13 14.09 22.66 24.97 13.97 21.84
SSIM 20.75 59.45 76.64 88.35 86.08 72.14 75.05 67.55 69.18 86.77
Img8 PSNR 8.56 17.30 18.74 16.78 26.03 14.78 19.41 23.02 11.32 22.75
SSIM 21.37 79.30 89.19 89.19 92.00 86.82 59.35 87.93 53.43 95.74
Img9 PSNR 10.38 16.05 14.82 18.56 22.70 18.23 21.62 19.20 13.91 21.51
SSIM 19.08 68.07 75.70 87.78 88.32 88.32 62.47 70.80 67.87 88.01
Img10 PSNR 10.43 15.82 14.95 20.89 18.04 15.33 20.38 17.12 13.64 22.80
SSIM 21.86 69.82 81.26 93.46 83.16 80.19 50.57 83.73 58.81 91.31
Img11 PSNR 10.58 14.66 14.89 18.06 21.91 15.97 16.69 18.70 17.75 19.86
SSIM 24.15 57.93 70.24 83.65 88.94 81.08 76.29 72.02 86.43 83.31
Img12 PSNR 10.90 16.69 14.17 18.64 25.52 19.39 17.51 21.27 13.74 20.54
SSIM 22.56 58.48 70.69 86.78 89.80 79.47 50.01 65.21 62.64 79.87
Img13 PSNR 11.87 15.57 14.03 20.33 16.95 21.76 17.54 23.34 14.89 22.94
SSIM 15.96 63.84 73.77 87.34 76.62 83.36 46.19 81.98 50.67 85.86
Img14 PSNR 9.93 14.44 18.84 17.28 17.10 16.68 14.87 17.49 20.39 18.85
SSIM 26.65 56.29 77.12 80.79 76.08 76.56 65.58 76.38 82.60 77.55
Img15 PSNR 9.30 14.86 17.67 16.29 16.05 15.99 12.36 15.42 16.25 18.55
SSIM 27.24 63.73 79.25 83.10 78.05 78.36 65.64 78.02 77.67 84.29
Ave. PSNR 9.95 15.54 17.64 17.83 18.39 16.56 15.91 17.97 15.85 20.21
SSIM 24.02 64.50 79.00 84.52 78.81 79.28 62.11 74.31 71.79 84.82
Three paired datasets Fivek [50], LSRW [56], and SYD [5] are tested to verify the effectiveness of the proposed method. We report average numerical image enhancement results of 600600600 images in Fivek, 965965965 images in SYD, and 505050 images in LSRW with PSNR, SSIM, NIQE, FADE, and SSEQ indexes in Table II. Besides, 151515 low-light images in the dataset RealBlur [57] with paired bright-light images are tested to demonstrate the validity of the proposed approach. Here we indicate these real low-light images as RealBlur15. We present the PSNR and SSIM results of dataset Realblur15 in Table III. The best and the second best values are marked as red and blue, respectively. From numerical results, the proposed method with the learned prior is better than the methods with traditional priors in most images. While comparing to the deep learning methods, some results fail to achieve the best one, most of our results still get the second best index. Our method achieves the best averaged overall performance among all the methods, which shows the effectiveness of our method on real low-light images.
(a) (a) Input (7.27/32.59)
(b) (b) RetinexNet (15.89/84.70)
(c) (c) GLADNet (20.16/79.30)
(d) (d) URetinexNet (19.07/83.12)
(e) (e) EFINet (19.43/89.01)
(f) (f) RQ (21.51/76.72)
(g) (g) SMG (16.35/84.87)
(h) (h) Bread (11.75/66.87)
(i) (i) Ours (21.65/90.12)
(j) (j) ExpertC
Figure 10: Image enhancement results with index PSNR (dB) and SSIM (%). (a) is the input low-light image from dataset FiveK; the enhanced results of (b) RetinexNet [19], (c) GLADNet [51], (d) URetinexNet [10], (e) EFINet [52], (f) RQ [53], (g) SMG [4], (h) Bread [54], and (i) Ours; (j) is the reference image renditions by expert C in FiveK dataset.
(a) (a) Input (10.35/40.65)
(b) (b) RetinexNet (14.64/78.98)
(c) (c) GLADNet (14.46/72.81)
(d) (d) URetinexNet (12.43/73.42)
(e) (e) EFINet (17.77/80.96)
(f) (f) RQ (16.01/70.50)
(g) (g) SMG (17.80/79.63)
(h) (h) Bread (13.73/61.83)
(i) (i) Ours (17.86/83.64)
(j) (j) ExpertC
Figure 11: Image enhancement results with index PSNR (dB) and SSIM (%). (a) is the input low-light image from dataset FiveK; the enhanced results of (b) RetinexNet [19], (c) GLADNet [51], (d) URetinexNet [10], (e) EFINet [52], (f) RQ [53], (g) SMG [4] (h) Bread [54], and (i) Ours; (j) is the reference image renditions by expert C in FiveK dataset.
(a) (a) Input (10.16/23.81)
(b) (b) RetiNet (15.25/72.64)
(c) (c) EFINet (16.14/86.83)
(d) (d) RQ (17.16/81.58)
(e) (e) LR3M (18.80/87.60)
(f) (f) QSPE (13.29/61.29)
(g) (g) DRE (14.83/81.71)
(h) (h) RWR (14.14/75.39)
(i) (i) Ours (19.01/88.26)
(j) (j) GT
Figure 12: Image enhancement results with indexes PSNR (dB) and SSIM (%). (a) is the observed low-light image Img1; the enhanced results of (b) RetinexNet [19], (c) EFINet [52], (d) RQ [53], (e) LR3M [14], (f) QSPE [55], (g) DRE [6], (h) RWR [18], and (i) Ours; (j) is the ground truth (GT).
Besides the numerical results, visual performance is presented in Figs. 10, 11. With the guidance of our learned edge extractor, the enhanced results in our (i) showed better structural preserving ability. While the SMG [4] is also guided by the edge module, some detailed information is not satisfactory. From both the numerical and visual results, the proposed method generates competitive performance. On the other hand, the visual results of Img1 from Realblur15 are presented in Fig 12. The real scene image with non-uniformed light and small spark lights is hard to handle. As shown in Fig. 12 (b), (c), and (e), we can see that the small spark lights affect the image enhancement results and make the lights abnormal. Compared to the paired ground-truth image (j), Fig. 12 (b), (d), (f), and (g) have a color bias due to the non-uniformed lights. Although the proposed approach generates the best results among the presented images, the color of (i) is still not exactly the same as the ground-truth image (j). Therefore, image enhancement for real-world images is still a challenging task.
V-E Non-reference quality assessment
In this subsection, five classical real-world non-reference datasets, DICM [58], LIME [59], MEF [60], Fusion [19], and NPE [61] are tested to verify the robustness of the proposed scheme. The state-of-the-art methods EFINet [52], Bread [54], RQ [53], LR3M [14], DRE [6], RMR [18] are compared. Due to the non-reference of the images, we conduct the non-reference metric ARISM, NIQE, and FADE as our index. In Table IV, the average numerical results of five datasets are listed. With the help of the learned edge extractor, the proposed method shows better performance compared to the approaches with the traditional priors. For most circumstances, the proposed method has a competitive result compared to the deep learning-based methods. Notably, our method achieved the highest ARISM score, which further demonstrates its effectiveness in enhancing real-world low-light images.
TABLE IV: Average image enhancement results on non-reference datasets DICM, LIME, MEF, NPE, and Fusion. We highlighted the BEST result in red and the SECOND BEST one in blue.
Datasets Index EFINet Bread RQ LR3M DRE RWR Ours
[…]
(a) (a) Input
(b) (b) EFINet (2.247)
(c) (c) Bread (2.535)
(d) (d) RQ (3.144)
(e) (e) LR3M (3.403)
(f) (f) DRE (2.144)
(g) (g) RWR (3.365)
(h) (h) Ours (2.139)
[…]
(a) (a) Input
(b) (b) EFINet (4.220)
(c) (c) Bread (5.132)
(d) (d) RQ (4.424)
(e) (e) LR3M (4.801)
(f) (f) DRE (4.586)
(g) (g) RWR (5.244)
(h) (h) Ours (3.519)
[…]
(a) (a) Input
(b) (b) EFINet (3.333)
(c) (c) Bread (3.779)
(d) (d) RQ (4.161)
(e) (e) LR3M (3.935)
(f) (f) DRE (3.597)
(g) (g) RWR (4.054)
(h) (h) Ours (3.224)
[…]
In Fig. 13, we present the image enhancement results of the image in dataset DICM. The input image (a) is quite simple to handle compared to other extreme dark images. However, the zoom part of results (c), (d), and (g) remain dark. Compared to other methods, our approach generates the enhanced image (h) with a normal light. An image from dataset LIME with edge feature is presented in Fig. 14. For better understanding, we zoom the texture part, which is a keyboard. Our result (h) has clear details of the keyboard in the enhanced image. From the zoom part of the input image (a), the keyboard is under the non-uniformed light, hence (b), (c), (e), and (f) are with the color bias. For (d) and (g), though the color remains, the details cannot be well preserved. In Fig. 15, we present the image enhancement results of the image in dataset MEF. Similar color bias happens in (d), (e), and (g). For better visuality, we zoom in on the mug in the image. Due to the help of the proposed edge network, our result (h) has a better enhancement performance while others still have low contrast. In conclusion, the proposed method achieved competitive results in enhancing real-world low-light images, as evidenced by both numerical and visual evaluations. The design of the learned edge extractor makes the proposed method particularly effective in handling texture images.
VI Conclusion
This paper proposed an edge-guided Retinex model with a novel inertial Bregman alternating linearized minimization algorithm for low-light image enhancement. With the help of deep learning, we trained the edge extractor with multi-scale residual blocks to obtain the edge regularizer. The learned edge extractor allows us to directly capture edge features from the low-light image, which helps the image enhancement. Additionally, we designed an efficient inertial Bregman alternating linearized minimization algorithm to tackle the proposed Retinex model. The effectiveness of the proposed algorithm was demonstrated through both theoretical and experimental analysis. To better illustrate the superiority of the proposed scheme, some state-of-the-art models on real-world datasets were compared. Based on the presented experimental results, the proposed method with learned prior and solved by the proposed algorithm achieves competitive performance in enhancing the low-light images. However, we acknowledge that our results still exhibit a color bias issue when dealing with non-uniform real-world low-light images. To address this, we plan to design an efficient model in the future to improve the performance of non-uniformed low-light image enhancement.